# Enumerating Finite Braid Group Orbits on $SL_{2}(\mathbb{C})$-Character Varieties

### v2

@@ 5 Primitive Reflection Groups
For each rank $r\in\{3,4\}$, we investigate the existence of nice $T$-tuples for $T=r$ and $T=r+1$ in the primitive reflection groups, denoted $G_{i}$ for $i=23,24,\dots,32$. For each group and each choice of $T$, we provide a table of the different types (up to inverse pairs), including the label for the type, the residues of the eigenvalues, as well as the nontrivial eigenvalues of the exemplar tuple under ‘Tuple Eigenvalues’. We also list the chosen exemplars $\mathbf{A}=[A_{1},A_{2},\dots,A_{T}]$ for each type. For each exemplar we compute the middle convolution with the suitable choice(s) of parameter $\lambda$ and this data is available in the appendices A, B, C. These tables also include, for each choice of eigenvalue $\lambda$, the residue of $\lambda$ in the ‘$\xi$’ column as well as the notation used in Magma to identify the root of unity under ‘$\lambda$’. We list the $2\times 2$ matrices composing $MC(\mathbf{A})=[M_{1},\dots,M_{T}]$ and record the order of the $GL_{2}(\mathbb{C})$ subgroup generated by the these matrices under ‘S. Size’. Note that a zero in this column means the subgroup is infinite.
@@
This is one reason why our braid group code optionally can return the list of signatures in each orbit. Note that this differs from the broader notion of equivalence in other literature such as Tykhyys’s classification [31]. Note that the work of Lisovy-Tykhyy [22] and Tykhyy [31] only consider the case where matrices in the tuple do not have a shared eigenvector. In our case, since nice tuples give an irreducible representation and middle convolution preserves irreducibility, we know all our tuples cannot share a common eigenvector. Thus, we expect to identify which tuples in these classifications come from the complex reflection groups via middle convolution. Details on what parameters we use to compare our orbits with these classifications is provided below at the start of Section 5.1 and 5.2.
@@ 5.1 Nice 3-Tuples in Rank 3 Groups
Nice tuples in this case will be tuples of 3 reflections generating the whole group such that their inverse product has some nontrivial eigenvalue of multiplicity 1. After middle convolution, our the induced tuple will correspond to $SL_{2}(\mathbb{C})$ local systems on the 4-punctured sphere and we can compare our braid group orbits to those in the literature. We compute the parameters $(\omega_{X},\omega_{Y},\omega_{Z},\omega_{4})$ in the notation of [22] and found that the $\omega$ parameters for each of our orbits match (up to permutation) with one listed in the work of Lisovy-Tykhyy. Since Tykhyy’s classification [31] further refines the original classification by Lisovy-Tykhyy [22], each tuple of $\omega$ parameters gives multiple orbits that Tykhyy identifies uniquely in terms of residues of the trace coordinates $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})$ and $(\sigma_{12},\sigma_{23},\sigma_{13},\sigma_{24})$ for some representative. In the nice 3-tuple case, using the $\omega$ parameters, we are able to easily identify how our orbit size relates the orbit size listed by Tykhyy for each $\mathcal{M}^{(4)}$ (in Tykhyy’s notation) and narrow down the possibilities for which of Tykhyy’s orbits our tuple corresponds to. In Tykhyy’s classification, tuples are considered equivalent up to: multiplying any two matrices by -1, complex conjugation of all elements of all matrices, cyclically permuting the matrices in the tuple, and up to taking inverses and reversing the order of the tuple [31]. We perform all these operations and compute the residue signatures for the matrices in each orbit to find one that matches the example provided in Tykhyy’s work. In most cases we found exact matches, and in other cases we provide the closest match, marking the differing coordinate with an asterisk (see Section 5.1.3 and 5.1.4).
Note that we only list this data for orbits that are considered different by our definition and these may end up being the same orbit in the Lisovy-Tykhyy or Tykhyy classifications ([22], [31]). For each orbit we compute, we provide the identifying label used in both the classifications, as well as our orbit size and the orbit size they provided (which is the same in both works).
@@ 5.1.1 $G_{23}$
The group $G_{23}$ of order 120 is also denoted $W(H_{3})$ with degrees $\deg(G_{23})=\{2,6,10\}$ [21, Table D.3]. All reflections in this group are conjugate with nontrivial eigenvalue -1. For nice 3-tuples in $G_{23}$, the computation returned three possible sets of eigenvalues for the inverse product.
@@
A $(-1,-1,-1)$ $\{\frac{1}{2},\frac{1}{10},\frac{9}{10}\}$
B $(-1,-1,-1)$ $\left\{\frac{1}{2},\frac{3}{10},\frac{7}{10}\right\}$
C $(-1,-1,-1)$ $\{\frac{1}{2},\frac{1}{6},\frac{5}{6}\}$
@@
The data for all three types for $G_{23}$ is compiled in Table A. For all three types, choosing a different exemplar would not have resulted in a new orbit. The two orbits of size 10 arising from Type A are the same, as are the two orbits of size 10 from Type B. However, these two orbits are distinct across the two types in the sense that their signatures did not show up in the others’ orbit. For Type C, the two orbits of size 18 are the same. Information about how these 6 distinct orbits relate to the other classifications is provided in Table 2. In this group we also find nice 4-tuples, details provided in Section 5.2.1
@@
$-q^{3}$ $(3-\sqrt{5},3-\sqrt{5},3-\sqrt{5},\frac{7\sqrt{5}-11}{2})$ 16 $(\frac{3}{5},\frac{2}{5},\frac{3}{5},\frac{2}{5})$ $(\frac{2}{5},\frac{2}{5},\frac{1}{3},\frac{3}{5})$ 57 10 10 $\langle 120,5\rangle$
@@
$-q^{4}$ $(3+\sqrt{5},3+\sqrt{5},3+\sqrt{5},-\frac{7\sqrt{5}+11}{2})$ 17 $(\frac{4}{5},\frac{4}{5},\frac{4}{5},\frac{4}{5})$ $(\frac{1}{5},0,\frac{1}{3},\frac{1}{3})$ 59 10 10 $\langle 120,5\rangle$
@@
-1 $(2,2,2,-1)$ 31 $(1,\frac{1}{3},0,0)$ $(\frac{1}{5},1,\frac{2}{5},\frac{2}{5})$ 99 18 72 0
$z^{5}$ $(2,2,2,-1)$ 31 $(\frac{2}{3},\frac{2}{3},\frac{2}{3},\frac{2}{3})$ $(0,\frac{1}{5},\frac{3}{5},\frac{3}{5})$ 96 18 18 $\langle 120,5\rangle$
@@ 5.1.2 $G_{24}$
The group $G_{24}$ of order 336 is also denoted $W(J_{3}(4))$ with degrees $\deg(G_{24})=\{4,6,14\}$ [21, Table D. 3]. All reflections in this group are conjugate with nontrivial eigenvalue -1. Looking at the degrees, we first expand the base field to be the cyclotomic field of degree 84. There is one inverse pair of nice 3-tuples in this group, so we only provide the details for one in Table 3.
@@
A $(-1,-1,-1)$ $\{\frac{1}{14},\frac{9}{14},\frac{11}{14}\}$
@@
[…]
The middle convolution data for $G_{24}$ is provided in Table A. These three distinct orbits are the only orbits from Type A and we checked that in all three cases, the inverse of the exemplars do not produce any new orbits. The comparison data for $G_{24}$ is provided in 4.
@@
$-z^{4}$ $(1,1,1,0)$ $8$ $(\frac{4}{7},\frac{4}{7},\frac{3}{7},\frac{1}{7})$ $(\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{2}{3})$ 28 7 28 0
$-z$ $(1,1,1,0)$ 8 $(\frac{1}{7},\frac{5}{7},\frac{1}{7},\frac{1}{7})$ $(\frac{1}{2},\frac{1}{2},\frac{1}{3},\frac{1}{2})$ 30 7 28 0
$-z^{2}$ $(1,1,1,0)$ 8 $(\frac{2}{7},\frac{2}{7},\frac{4}{7},\frac{2}{7})$ $(\frac{1}{2},\frac{1}{2},\frac{1}{3},\frac{1}{2})$ 29 7 28 0
@@ 5.1.3 $G_{25}$
The group $G_{25}$ of order 648 is also denoted $W(L_{3})$ with degrees $\deg(G_{25})=\{6,9,12\}$ [21, Table D.3]. $G_{25}$ has two conjugacy classes, consisting of matrices with nontrivial eigenvalue $w$ or $w^{2}$ respectively, for $w=\zeta_{3}$. Expanding the base field to be the cyclotomic field of order 36, we get 8 possible types, which actually split into 4 pairs of inverses. We only list one set of inverse eigenvalues for each pair in Table 5.
@@ D $(w,w,w)$$\{\frac{1}{3},\frac{5}{6},\frac{5}{6}\}$
The details for the middle convolution of each type is provided in Table 27. For all three types, different exemplars do not produce new orbits nor do their inverses. For Type A, the two orbits of size 12 are distinct, as are all three orbits of size 36 from Type B. For Type C, we had to extend to a field containing $\zeta_{72}$ to construct the character. The two orbits of size 4 from Type C are the same. Including the one from Type D, we get 8 distinct orbits from $G_{25}$. The comparison data is provided in Table 6. Comparing with the Lisovy-Tykhyy paper, all the $\omega$ parameters seemed to correspond to the family of orbits listed in Lemma 39 in [22], so we have provided the specific values for the variables that work in each case. When comparing with Tykhyy’s classification, there was only one orbit (from Type B) that we could not find an exact match for. We listed the closest match, which seemed to be Orbit 5 in [31] and distinguish the different coordinate with an asterisk. In this group we also find nice 4-tuples, details provided in Section 5.2.2
@@
$-w$ $(0,0,3,-2)$ Type II, $X^{\prime}=2,X^{\prime\prime}=1$ $(\frac{1}{6},\frac{1}{2},\frac{1}{6},\frac{1}{2})$ $(\frac{1}{2},\frac{1}{2},\frac{1}{3},0)$ 3 2 12 $\langle 24,4\rangle$
$-w^{2}$ $(0,0,3,-2)$ Type II, $X^{\prime}=2,X^{\prime\prime}=1$ $(\frac{1}{6},\frac{1}{2},\frac{1}{6},\frac{1}{2})$ $(\frac{1}{2},\frac{1}{2},\frac{1}{3},0)$ 3 2 12 $\langle 24,4\rangle$
$w^{2}$ $(0,0,3,-2)$ Type II, $X^{\prime}=2,X^{\prime\prime}=1$ $(1,\frac{1}{3},1\frac{2}{3})$ $(\frac{1}{2},\frac{1}{2},0,\frac{1}{3})$ 4 2 24 0
@@
$-z^{10}$ (2, -1, -1, -1) Type III, $\omega=-1$ $(\frac{8}{9},\frac{4}{9},\frac{4}{9},\frac{2}{3})$ $(\frac{1}{2},\frac{1}{3},\frac{1}{2},\frac{2}{3}^{*})$ $5^{*}$ 3 36 0
@@
$-z^{9}+z^{3}$ $(4,4,4,-8)$ Type IV, $\omega=4$ $(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})$ $(\frac{1}{3},\frac{1}{3},\frac{1}{3},0)$ 6 4 4 $\langle 48,28\rangle$
$-z^{6}$ $(4,4,4,-8)$ Type IV, $\omega=4$ $(0,0,0,\frac{1}{2})$ $(\frac{1}{3},\frac{1}{3},\frac{1}{3},0)$ 7 4 4 0
@@
$w$ $(3,3,3,-5)$ Type I, $X=Y=Z=1$ $(\frac{2}{3},\frac{2}{3},0,\frac{1}{3})$ $(\frac{1}{3},\frac{2}{3},\frac{2}{3},\frac{2}{3})$ 2 1 4 $\langle 24,3\rangle$
Table 6: Here $w=\zeta_{3},z=\zeta_{36}$. The correct signature for Tykhyy’s orbit $5$ would have $3(\frac{4}{9})\equiv\frac{1}{3}$ as $\sigma_{24}$ but such a signature was not in the orbit, we have provided the closest match. The subgroups $\langle 24,4\rangle,\langle 48,28\rangle,\langle 24,3\rangle$ correspond to the dicyclic group of order 24 (binary dihedral group), binary octahedral group, and $SL(2,3)$ respectively.

### v3

@@ 5 Primitive Reflection Groups
For each rank $r\in\{3,4\}$, we investigate the existence of nice $T$-tuples for $T=r$ and $T=r+1$ in the primitive reflection groups, denoted $G_{i}$ for $i=23,24,\dots,32$. For each group and each choice of $T$, we provide a table of the different types (up to inverse pairs), including the label for the type, the residues of the eigenvalues, as well as the nontrivial eigenvalues of the exemplar tuple under ‘Tuple Eigenvalues’. We also list the chosen exemplars $\mathbf{A}=[A_{1},A_{2},\dots,A_{T}]$ for each type. For each exemplar we compute the middle convolution with the suitable choice(s) of parameter $\lambda$ and this data is available in the appendices A, B, C. These tables also include, for each choice of eigenvalue $\lambda$, the residue of $\lambda$ in the $\xi$’ column as well as the notation used in Magma to identify the root of unity under $\lambda$’. We list the $2\times 2$ matrices composing $MC(\mathbf{A})=[M_{1},\dots,M_{T}]$ and record the order of the $GL_{2}(\mathbb{C})$ subgroup generated by the these matrices under ‘S. Size’. Note that a zero in this column means the subgroup is infinite.
@@
This is one reason why our braid group code optionally can return the list of signatures in each orbit. Note that this differs from the broader notion of equivalence in other literature such as Tykhyys’s classification [Tyk]. Note that the work of Lisovyy-Tykhyy [T-L] and Tykhyy [Tyk] only consider the case where matrices in the tuple do not have a shared eigenvector. In our case, since nice tuples give an irreducible representation and middle convolution preserves irreducibility, we know all our tuples cannot share a common eigenvector. Thus, we expect to identify which tuples in these classifications come from the complex reflection groups via middle convolution. Details on what parameters we use to compare our orbits with these classifications is provided below at the start of Section 5.1 and 5.2.
@@ 5.1 Nice 3-Tuples in Rank 3 Groups
Nice tuples in this case will be tuples of 3 reflections generating the whole group such that their inverse product has some nontrivial eigenvalue of multiplicity 1. After middle convolution, the induced tuple will correspond to $SL_{2}(\mathbb{C})$ local systems on the 4-punctured sphere and we can compare our braid group orbits to those in the literature. We compute the parameters $\omega_{X},Y,Z,4=(\omega_{X},\omega_{Y},\omega_{Z},4-\omega_{4})$ in the notation of [T-L, Equations 11-12] and found that the $\omega$ parameters for each of our orbits match (up to permutation) with one listed in the work of Lisovyy-Tykhyy. Since Tykhyy’s classification [Tyk] further refines the original classification by Lisovyy-Tykhyy [T-L], each tuple of $\omega$ parameters gives multiple orbits that Tykhyy identifies uniquely in terms of residues of the trace coordinates $\theta_{1,2,3,4}=(\theta_{1},\theta_{2},\theta_{3},\theta_{4})$ and $\sigma_{12,23,13,24}=(\sigma_{12},\sigma_{23},\sigma_{13},\sigma_{24})$ for some representative. In the nice 3-tuple case, using the $\omega$ parameters, we are able to easily identify how our orbit size relates the orbit size listed by Tykhyy for each $\mathcal{M}^{(4)}$ (in Tykhyy’s notation) and narrow down the possibilities for which of Tykhyy’s orbits our tuple corresponds to. In Tykhyy’s classification, tuples are considered equivalent up to: multiplying any two matrices by -1, complex conjugation of all elements of all matrices, cyclically permuting the matrices in the tuple, and up to taking inverses and reversing the order of the tuple [Tyk]. We perform all these operations and compute the residue signatures based on Tykhyy’s definition [Tyk, Section 2, (1)-(3)] for the matrices in each orbit to find one that matches an example provided in Table 1 of Tykhyy’s work. In most cases we find an exact match, with a few exceptions corresponding to Orbit 5 in Tykhyy’s Table 1. For those tuples, we mark the differing coordinate with an asterisk (see Section 5.1.3 and 5.1.4), noting that they still agree with Tykhyy’s list as the $\theta$ and $\sigma$ coordinates are defined modulo 2 and up to a sign in Tykhyy’s work [Tyk, Section 2, (3)].
Note that we only list this data for orbits that are considered distinct by our Definition 5.0.1 and these may end up being the same orbit in the Lisovyy-Tykhyy or Tykhyy classifications ([T-L], [Tyk]). For each orbit we compute, we provide the identifying label used in both the classifications, as well as our orbit size and the orbit size they provided (which is the same in both works).
@@ 5.1.1 $G_{23}$
The group $G_{23}$ of order 120 is also denoted $W(H_{3})$ with degrees $\deg(G_{23})=\{2,6,10\}$ [ld2009, Table D.3]. All reflections in this group are conjugate with nontrivial eigenvalue -1. For nice 3-tuples in $G_{23}$, the computation returned three possible sets of eigenvalues for the inverse product.
@@
A (1,1,1)(-1,-1,-1) $\{\frac{1}{2},\frac{1}{10},\frac{9}{10}\}$
B (1,1,1)(-1,-1,-1) $\left\{\frac{1}{2},\frac{3}{10},\frac{7}{10}\right\}$
C (1,1,1)(-1,-1,-1) $\{\frac{1}{2},\frac{1}{6},\frac{5}{6}\}$
@@
The data for all three types for $G_{23}$ is compiled in Table 31. For all three types, choosing a different exemplar would not have resulted in a new orbit. The two orbits of size 10 arising from Type A are the same, as are the two orbits of size 10 from Type B. However, these two orbits are distinct across the two types in the sense that their signatures did not show up in the others’ orbit. For Type C, the two orbits of size 18 are the same. Information about how these 6 distinct orbits relate to the other classifications is provided in Table 6. In this group we also find nice 4-tuples, details provided in Section 5.2.1
@@
$q^{3}$ $(3-\sqrt{5},3-\sqrt{5},3-\sqrt{5},\frac{7\sqrt{5}-11}{2})$ 16 $(\frac{3}{5},\frac{2}{5},\frac{3}{5},\frac{2}{5})$ $(\frac{2}{5},\frac{2}{5},\frac{1}{3},\frac{3}{5})$ 57 10 10 $\langle 120,5\rangle$
@@
$q^{4}$ $(3+\sqrt{5},3+\sqrt{5},3+\sqrt{5},-\frac{7\sqrt{5}+11}{2})$ 17 $(\frac{4}{5},\frac{4}{5},\frac{4}{5},\frac{4}{5})$ $(\frac{1}{5},0,\frac{1}{3},\frac{1}{3})$ 59 10 10 $\langle 120,5\rangle$
@@
-1 (2,2,2,1)(2,2,2,-1) 31 $(1,\frac{1}{3},0,0)$ $(\frac{1}{5},1,\frac{2}{5},\frac{2}{5})$ 99 18 72 0
$z^{5}$ (2,2,2,1)(2,2,2,-1) 31 $(\frac{2}{3},\frac{2}{3},\frac{2}{3},\frac{2}{3})$ $(0,\frac{1}{5},\frac{3}{5},\frac{3}{5})$ 96 18 18 $\langle 120,5\rangle$
@@ 5.1.2 $G_{24}$
The group $G_{24}$ of order 336 is also denoted $W(J_{3}(4))$ with degrees $\deg(G_{24})=\{4,6,14\}$ [ld2009, Table D. 3]. All reflections in this group are conjugate with nontrivial eigenvalue -1. Looking at the degrees, we first expand the base field to be the cyclotomic field of degree 84. There is one inverse pair of nice 3-tuples in this group, so we only provide the details for one in Table 7.
@@
A (1,1,1)(-1,-1,-1) $\{\frac{1}{14},\frac{9}{14},\frac{11}{14}\}$
@@
$A_{1}=\begin{pmatrix}1&-z^{4}-z^{2}-z&-z^{4}-z^{2}-z\\ \>0&0&-1\\ \>0&-1&0\end{pmatrix}\quad A_{2}=\begin{pmatrix}1&1&0\\ \>0&-1&0\\ \>0&1&1\end{pmatrix}\quad A_{3}=\begin{pmatrix}-1&0&0\\ \>1&1&0\\ \>z^{4}+z^{2}+z&0&1\end{pmatrix}$
The middle convolution data for $G_{24}$ is provided in Table 32. These three distinct orbits are the only orbits from Type A and we checked that in all three cases, the inverse of the exemplars do not produce any new orbits. The comparison data for $G_{24}$ is provided in 8.
@@
$z^{4}$ $(1,1,1,0)$ 88 $(\frac{4}{7},\frac{4}{7},\frac{3}{7},\frac{1}{7})$ $(\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{2}{3})$ 28 7 28 0
z-z $(1,1,1,0)$ 8 $(\frac{1}{7},\frac{5}{7},\frac{1}{7},\frac{1}{7})$ $(\frac{1}{2},\frac{1}{2},\frac{1}{3},\frac{1}{2})$ 30 7 28 0
$z^{2}$ $(1,1,1,0)$ 8 $(\frac{2}{7},\frac{2}{7},\frac{4}{7},\frac{2}{7})$ $(\frac{1}{2},\frac{1}{2},\frac{1}{3},\frac{1}{2})$ 29 7 28 0
@@ 5.1.3 $G_{25}$
The group $G_{25}$ of order 648 is also denoted $W(L_{3})$ with degrees $\deg(G_{25})=\{6,9,12\}$ [ld2009, Table D.3]. $G_{25}$ has two conjugacy classes, consisting of matrices with nontrivial eigenvalue $w$ or $w^{2}$ respectively, for $w=\zeta_{3}$. Expanding the base field to be the cyclotomic field of order 36, we get 8 possible types, which actually split into 4 pairs of inverses. We only list one set of inverse eigenvalues for each pair in Table 9.
@@ D $(w,w,w)$ $\{\frac{1}{3},\frac{5}{6},\frac{5}{6}\}$
The details for the middle convolution of each type is provided in Table 33. For all three types, different exemplars do not produce new orbits nor do their inverses. For Type A, the two orbits of size 12 are distinct, as are all three orbits of size 36 from Type B. For Type C, we had to extend to a field containing $\zeta_{72}$ to construct the character. The two orbits of size 4 from Type C are the same. Including the one from Type D, we get 8 distinct orbits from $G_{25}$. The comparison data is provided in Table 10. Comparing with the Lisovyy-Tykhyy paper, all the $\omega$ parameters seemed to correspond to the family of orbits listed in Lemma 39 in [T-L], so we have provided the specific values for the variables that work in each case. In this group we also find nice 4-tuples, details provided in Section 5.2.2
@@
w-w (0,0,3,2)(0,0,3,-2) Type II, $X^{\prime}=2,X^{\prime\prime}=1$ $(\frac{1}{6},\frac{1}{2},\frac{1}{6},\frac{1}{2})$ $(\frac{1}{2},\frac{1}{2},\frac{1}{3},0)$ 3 2 12 $\langle 24,4\rangle$
$w^{2}$ (0,0,3,2)(0,0,3,-2) Type II, $X^{\prime}=2,X^{\prime\prime}=1$ $(\frac{1}{6},\frac{1}{2},\frac{1}{6},\frac{1}{2})$ $(\frac{1}{2},\frac{1}{2},\frac{1}{3},0)$ 3 2 12 $\langle 24,4\rangle$
$w^{2}$ (0,0,3,2)(0,0,3,-2) Type II, $X^{\prime}=2,X^{\prime\prime}=1$ $(1,\frac{1}{3},1,\frac{2}{3})$ $(\frac{1}{2},\frac{1}{2},0,\frac{1}{3})$ 4 2 24 0
@@
$z^{10}$ (2, -1, -1, -1) Type III, $\omega=-1$ $(\frac{8}{9},\frac{4}{9},\frac{4}{9},\frac{2}{3})$ $(\frac{1}{2},\frac{1}{3},\frac{1}{2},\frac{2}{3}^{*})$ $5^{*}$ 3 36 0
@@
$z^{9}+z^{3}$ (4,4,4,8)(4,4,4,-8) Type IV, $\omega=4$ $(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})$ $(\frac{1}{3},\frac{1}{3},\frac{1}{3},0)$ 6 4 4 $\langle 48,28\rangle$
$z^{6}$ (4,4,4,8)(4,4,4,-8) Type IV, $\omega=4$ $(0,0,0,\frac{1}{2})$ $(\frac{1}{3},\frac{1}{3},\frac{1}{3},0)$ 7 4 4 0
@@
$w$ (3,3,3,5)(3,3,3,-5) Type I, $X=Y=Z=1$ $(\frac{2}{3},\frac{2}{3},0,\frac{1}{3})$ $(\frac{1}{3},\frac{2}{3},\frac{2}{3},\frac{2}{3})$ 2 1 4 $\langle 24,3\rangle$
Table 10: Here $w=\zeta_{3},z=\zeta_{36}$. The subgroups $\langle 24,4\rangle,\langle 48,28\rangle,\langle 24,3\rangle$ correspond to the dicyclic group of order 24 (binary dihedral group), binary octahedral group, and $SL(2,3)$ respectively. The signature for Tykhyy’s orbit $5$ is $\theta_{1,2,3,4}=(2x,x,x,\frac{2}{3})$ and $\sigma_{12,23,13,24}=(\frac{1}{2},\frac{1}{3},\frac{1}{2},3x)$ but such a signature was not in the orbit so we provide the closest match. The coordinate that differs from the formula is marked with an asterisk, but note that as $\sigma_{24}$ is only defined modulo 2 and up to sign, we believe this still corresponds to Orbit 5 of Tykhyy’s Table 1.